OpenStudy (anonymous):

Show that the equation has a root in the given interval: 3x^3-2x^2+5x+4=0 [-1,0] The book I have explains nothing...

OpenStudy (amistre64):

We would have to factor it into is baser parts; like 12 factors in 4 and 3

OpenStudy (amistre64):

we can try to find a way to group it into parts that would work; but thats as much trial and error as my way :)

OpenStudy (amistre64):

One method is to solve the equation for -1 and then for 0 and if there is a switch in signs; then it crosses the axis

OpenStudy (anonymous):

who hs gmail

OpenStudy (anonymous):

go to that

myininaya (myininaya):

do f(-1) and do f(0) if one is positive and one is negative then there is root on the interval -1..0

OpenStudy (amistre64):

like this example

OpenStudy (anonymous):

who `has gmail

OpenStudy (anonymous):

http://uk.answers.yahoo.com/question/index?qid=20110324085500AAVlcai - This link will have the answer.

OpenStudy (anonymous):

Okay, i'm with you

OpenStudy (anonymous):

Use the Intermediate Value Theorem to check for the existence of zeros (like myininaya said). If f(x) is continuous on [a, b] and if f(a) and f (b) are nonzero and have opposite signs, then f(x) has a zero in (a, b).

OpenStudy (amistre64):

3(-1)^3-2(-1)^2+5(-1)+4 = ?? -3 -2 -5 +4 = -6 3(0)^3-2(0)^2+5(0)+4 = +4 Since -6 and 4 differ by "sign" the function has to cross the xaxis at some point in the interval between [-1,0]

OpenStudy (anonymous):

Wow, I sure wish the book explained this stuff.

OpenStudy (anonymous):

ammmmmmmmmmmmmmmmmmmmmmmmmmistrrrrrrrrrrrrrrrrre6444444444444444444444444444444

OpenStudy (amistre64):

mary; please do not disrupt other peoples learning environment. If you do not have anything to contribute to the question at hand, please dont comment... thank you.

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

email it to me, because it doesnt come out on my end....

OpenStudy (anonymous):

whats your gmail

OpenStudy (amistre64):

Jenn: Does that explanation make sense for you? If you start on one side of the road and end up on the other side; youd have to have crossed the road at some point right?

OpenStudy (amistre64):

same as it was last time :) tony031172@gmail.com

OpenStudy (anonymous):

Yes, it makes more sense. This one is not as bad as the other one you helped me with.

OpenStudy (anonymous):

plz check your mail now

OpenStudy (amistre64):

lol...not as many chances to mess it up ;)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Too funny..i have several more of those as well, so i'll try it

OpenStudy (anonymous):

If I do them, will you let me know if the answer are right? Again,please..i can post them again too if you'd like

OpenStudy (amistre64):

mary; the assassination of lincoln again? no thnx :)

OpenStudy (amistre64):

We all can check ;)

OpenStudy (anonymous):

Great...thank you so very much.

OpenStudy (amistre64):

there no I in team, but there is an atem..and meat... and ate, tea eat, met, and other useful yet trivial words :)

OpenStudy (anonymous):

i need you to help me on google doc

OpenStudy (anonymous):

i needdd your hep

OpenStudy (amistre64):

mary: helps is the operative word there. I am good at math, lincolns assassination, while it was a long time ago, still happened a few years before me :)

OpenStudy (anonymous):

but plz go to gmail

OpenStudy (amistre64):

its blank....

OpenStudy (anonymous):

i nived you to my google doc

OpenStudy (anonymous):

amistre why you left

OpenStudy (anonymous):

so mean ou can help me an dget this done